OpenStudy (anonymous):

divide (y^4-81)/ (y-3) @SalvadorV

OpenStudy (anonymous):

that one is a little tricky, do you know how to divide functions?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

Factorise (y^4-81) and then cancel out the common factor i.e,(y-3).

OpenStudy (anonymous):

what you can see there is that you have Y^4+3^4, (3^4=81)

OpenStudy (anonymous):

sorry y^4-3^4

OpenStudy (anonymous):

@Onlymaths can you draw that out for me @SalvadorV yes

OpenStudy (anonymous):

\[y^4-3^4\]

OpenStudy (anonymous):

@SalvadorV where did you get 3^4 =81 from and why

OpenStudy (anonymous):

3^4= 3*3*3*3= 9*3*3=27*3=81

OpenStudy (anonymous):

y^4-3^4=(y^2-9)(y^2+9)

OpenStudy (anonymous):

so, we know that \[(a+b)(a-b)=a^2+ab-ab-b^2=a^2-b^2\]

OpenStudy (anonymous):

okay yes

OpenStudy (anonymous):

so then on the same logic \[(a^2-b^2)(a^2+b^2)=a^4+a^2b^2-a^2b^2-b^4=a^4-b^4\]

OpenStudy (anonymous):

do you see why?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

it is the same thing, just that they are now squared

OpenStudy (anonymous):

ookay yea go on

OpenStudy (anonymous):

so we have \[a^4-b^4=(a^2-b^2)(a^2+b^2)=((a+b)(a-b))(a^2+b^2)\]

OpenStudy (anonymous):

yea what is next

OpenStudy (anonymous):

you see, that is the problem we are looking for

OpenStudy (anonymous):

your answer is there, can you see how?

OpenStudy (anonymous):

umm no can you break it down more draw it please

OpenStudy (anonymous):

@SalvadorV

OpenStudy (anonymous):

say we have our\[a^4-b^4=((a-b)(a+b))(a^2+b^2)\] but I say a=y and b=3 then it would be \[y^4-3^4=((y-3)(y+3))(y^2+3^2)\]

OpenStudy (anonymous):

can you add the numbers in please

OpenStudy (anonymous):

so, can you see now, what we want to do is \[((y-3)(y+3)(y^2+3^2))/(y-3)\]

OpenStudy (anonymous):

it is \[y^4-81=(y-3)(y+3)(y^2-9)\]

OpenStudy (anonymous):

do you know what you have to do here?